# Understanding Before Recommendation: Semantic Aspect-Aware Review Exploitation via Large Language Models

### v1

@@ 3.1. Preliminary
Before describing our model, we would like to introduce the problem setting first. A dataset with an interaction matrix $\bm{R}^{N\times M}$ of a set of users $\mathcal{U}$ and a set of items $\mathcal{I}$ is involved, where $N$ and $M$ are the numbers of users and items, respectively. The matrix $\bm{R}$ records the interaction history between users and items. A nonzero entry $r_{ui}\in\bm{R}$ indicates that user $u\in\mathcal{U}$ has interacted with item ${i}\in\mathcal{I}$ before; otherwise, the entry is zero. For each user-item pair $(u,i)$, a review $v$ is associated with them. The review contains the user’s sentiment on different aspects of an item. We employ LLM to extract a semantic aspect set $\mathcal{A}$, which includes all the aspects mentioned in the reviews. Specifically, $\mathcal{A}_{v}=\{a_{v}^{1},\cdots,a_{v}^{N_{v}}\}$ is the set of semantic aspects extracted from review $v$, where $N_{v}$ is the number of extracted aspects. In this work, our goal is to explore the potential of the extracted semantic aspects from raw reviews by LLMs to enhance recommendation accuracy and interpretability.
@@
To achieve the goal, we (1) design a chain-based prompting approach, which can effectively extract semantic aspect-aware reviews automatically; (2) introduce a semantic aspect-based GCN model for recommendation. We use $\{\bm{R^{1}},\bm{R^{2}},\cdots,\bm{R^{A}}\}$ to denote the semantic aspect-based interaction matrices. $A$ is the number of aspects. $\bm{R^{a}}$ is the interaction matrix of the $a$-th aspect.
Figure 2. Chain-based Prompting Approach.
3.1.2. Semantic Aspect-based Graph Construction.
@@
In the next, we detail our method for semantic aspect-aware review extraction.
3.2. Semantic Aspect-Aware Review Extraction
In this section, we conduct sentiment analysis on user reviews using LLM to discern user-item interactions on various semantic aspects. As illustrated in Fig 2, we introduce a chain-based prompting approach by combining the semantic aspects and semantic aspect-aware review extraction.
3.2.1. Semantic Aspects Extraction
The semantic aspects vary depending on the recommendation scenario. For instance, when buying dresses, style is crucial, but for office products, ease of use is the main concern. Empirically, it is challenging to extract the high-quality unseen aspects and reviews simultaneously. Therefore, we first extract semantic aspects from all reviews of each dataset. The semantic aspects extraction prompt is designed as follows:
As shown in Fig. 1, LLM generates a semantic aspect set and its corresponding analysis for each aspect. Note that, we provide little information about the semantic aspect in our prompt, leading to low-quality analysis. By iteratively analyzing each review using LLM, we can obtain all semantic aspects within this recommendation domain by aggregating all the semantic aspect sets of all reviews.
3.2.2. Semantic Aspect-Aware Review Extraction
Based on the extracting semantic aspects, we further employ LLM to generate high-quality semantic aspect-aware reviews. By focusing on specific aspects, high-quality semantic aspect-aware reviews can be obtained. For instance, on the Baby Dataset, we get a semantic aspect set containing Quality, Comfort, Durability, Design, Functionality, Ease of use, Price, and Size. To accurately identify reviews to different aspects, we adapt the prompt to the semantic aspects of the recommendation scenario:
  • Prompt 2: A person bought a product and commented that ¡review¿. Tell me from which perspectives the customer gave this review, e.g., Quality, Comfort, Durability, Design, Functionality, Ease of Use, Price, and Size. Answer point by point.
The LLM identifies the semantic aspect-aware review for the aspects we input. Then, we identify the semantic aspect-aware interactions between the user and the item based on the generated semantic-aware reviews. For example, if the LLM generates the output: “The customer mentioned that the bins were not durable and that the fabric disintegrated when burned with a candle.”, it indicates an interaction on Functionality. On the other hand, if the LLM generates the output: “The customer did not mention anything about the ease of use of the bins.”, which means that there is no interaction on this aspect.
3.3. Semantic Aspect-Based Graph Convolution Network
Figure 3. Overview of our SAGCN model.
3.3.1. Overview
Based on the extracted semantic-aware user-item interactions for semantic aspects, we seek to model users’ preferences on various aspects with GCN-based techniques. The structure of our SAGCN is illustrated in Figure 1, and it is composed of three key components: 1) embedding initialization, which initialized node (users and items) embeddings for the subsequent learning in each aspect-specific interaction graph; 2) embedding propagation, in which LightGCN (He et al., 2020) is adopted to learn the user and item embeddings by performing graph convolution on each semantic aspect-based graph; 3) embedding generation, which aggregates the embeddings learned from each aspect-specific graph and combines the embeddings of different aspects.
@@ 3.3.2. Embedding Initialization
In our model, the vector of users and items is composed of embeddings from A aspects. Following existing GCN-based recommendation models (Wang et al., 2019b; He et al., 2020; Liu et al., 2021), we initialize the embedding vectors for each aspect. As an example, the ID embeddings of user $u\in\mathcal{U}$ for different aspects are initialized as:
@@
where $\bm{e_{u}^{(0)}}\in\mathds{R}^{d\times A}$ is ID embedding to capture intrinsic characteristics of user $u$; $\bm{e_{u}^{(a,0)}}\in\mathds{R}^{{d}}$ is $u$’s representation in aspect $a$. Analogously, $\bm{e_{i}}^{(0)}=\{\bm{e_{i}^{(1,0)},e_{i}^{(2,0)},\cdots,e_{i}^{(A,0)}}\}$ is defined as the ID embeddings of item $i\in\mathcal{I}$. The embeddings of the user ID and item ID for each aspect are randomly initialized in our model.
@@ 3.3.3. Embedding Propagation
To exploit the high-order connectives on each aspect to learn user and item representations, we propagate embeddings on multiple semantic aspect-based graphs. Take the semantic aspect-specific graph $\mathcal{G}_{a}$ as an example, let $\bm{e_{u}^{(a,k)}}$ and $\bm{e_{i}^{(a,k)}}$ denote the embedding of user $u$ and item $i$ after $k$ layers graph convolution, where $k\geq 0$, the embedding propagation is formulated as:
@@ 3.3.5. Model Training
The top-$n$ recommendation aims to recommend a set of $n$ top-ranked items matching the target user’s preference. Similar to previous works (Liu et al., 2021), we adopt the pairwise learning method for optimization. To train the proposed model, we construct a triplet of $\{u,i^{+},i^{-}\}$, with a positive item (an observed interaction between $u$ and $i^{+}$) and a negative item (an unobserved interaction between $u$ and $i^{-}$ [^1] Note that pairwise learning assumes that a positive item (i.e., $i^{+}$) should rank higher than a negative item (i.e., $i^{-}$).. The objective function is formulated as follows:
@@
where $\mathcal{O}=\{(u,i^{+},i^{-})|(u,i^{+})\in\mathcal{R^{+}},(u,i^{-})\in\mathcal{R^{-}}\}$ denotes the training set; $\mathcal{R^{+}}$ indicates the observed interactions between user $u$ and $i^{+}$ in the training dataset, and $\mathcal{R^{-}}$ is the sampled unobserved interaction set. $\lambda$ and $\Theta$ represent the regularization weight and the parameters of the model, respectively. The $L_{2}$ regularization is used to prevent overfitting. Besides, we adopt the mini-batch Adam (Kingma and Ba, 2015) to optimize the model and update its parameters.
@@ 4.1. Experimental setup
Table 2. Performance of our method and the competitors over three datasets.
Datasets Office Baby Clothing
k=10 k=20 k=10 k=20 k=10 k=20
Metrics Recall NDCG Recall NDCG Recall NDCG Recall NDCG Recall NDCG Recall NDCG
NeuMF 5.14 3.89 8.12 5.21 3.11 2.11 4.85 2.69 0.94 0.54 1.50 0.71
GCMC 6.72 5.27 10.27 6.79 4.55 2.99 7.24 3.89 3.17 1.86 4.86 2.35
LightGCN 9.87 6.04 14.47 7.43 5.94 3.30 9.25 4.20 4.45 2.43 6.44 2.95
DGCF 9.95 6.25 14.37 7.52 5.90 3.27 9.20 4.14 4.67 2.68 6.91 3.32
IMP-GCN 10.11 6.36 14.47 7.71 6.24 3.49 9.56 4.38 4.80 2.76 7.11 3.40
NCL 10.07 6.30 14.40 7.65 6.15 3.42 9.45 4.30 4.76 2.74 7.10 3.37
MMGCN 5.74 3.42 9.39 4.54 3.95 2.17 6.46 2.85 2.42 1.29 3.76 1.64
GRCN 10.38 6.34 15.33 7.81 5.57 3.03 8.49 3.83 4.47 2.35 6.70 2.94
LATTICE 10.00 6.09 14.99 7.57 6.06 3.40 9.29 4.27 5.03 2.79 7.28 3.37
BM3 9.80 6.09 14.02 7.36 6.44 3.65 9.52 4.48 5.28 2.94 7.75 3.58
SAGCN 11.71* 7.34* 16.71* 8.84* 7.35* 4.23* 10.56* 5.09* 6.07* 3.58* 8.44* 4.20*
Imporv. 12.81% 15.77% 9.00% 13.19% 14.13% 15.89% 10.46% 13.61% 14.96% 21.77% 8.90% 17.32%

### v2

@@ 3.1. Preliminary
To provide the details for our approach, let’s first outline the problem setting. We are working with a dataset that consists of an interaction matrix denoted as $\bm{R}^{N\times M}$, where $\mathcal{U}$ represents the set of users, and $\mathcal{I}$ represents the set of items. $N$ and $M$ represent the total number of users and items, respectively. The matrix $\bm{R}$ records the interaction history between users and items. A nonzero entry $r_{ui}\in\bm{R}$ indicates that user $u\in\mathcal{U}$ has interacted with item ${i}\in\mathcal{I}$ before; otherwise, the entry is zero. For each user-item pair $(u,i)$, a review $v$ is associated with them. In this work, our goal is to explore the potential of reviews to enhance recommendation accuracy and interpretability.
@@
To achieve this goal, we propose an approach that integrates LLMs with conventional recommendation techniques. Specifically, we (1) design a chain-based prompting approach, which can effectively extract semantic aspect-aware reviews without clear instructions; (2) introduce a semantic aspect-based GCN model for recommendation. We use $\{\bm{R^{1}},\bm{R^{2}},\cdots,\bm{R^{A}}\}$ to denote the semantic aspect-based interaction matrices. $A$ is the number of aspects. $\bm{R^{a}}$ is the interaction matrix of the $a$-th aspect.
3.2. Overview
Recommendation systems learn user and item representations by exploiting user-item interactions, such as clicks and reviews. In this way, users’ implicit preferences are encoded into latent vectors. Notably, user reviews offer explicit intents of a user towards an item. For example, the review might mention various semantic aspects such as functionality, durability, and ease of use. Each of these aspects carries a specific meaning related to user experience. Recognizing these explicit intentions leads to a deeper understanding of user preferences. More recently, LLMs have emerged as powerful tools for sentiment analysis (Sileo et al., 2022), capable of extracting semantic aspects from user reviews. Therefore, to enhance the accuracy of recommendation systems, a desired approach should well seamlessly combine the advanced capabilities of LLMs with conventional recommendation techniques.
Figure 2. Overview of Our Approach. Our approach consists of two components: sentiment analysis and representation learning. The former leverages the LLM with a Chain-based Prompting Strategy (short for CPS) to extract the semantic aspect-aware interactions from user reviews. The latter is dedicated to learning user and item representations based on these interactions, employing the Semantic Aspect-aware Graph Convolutional Network (short for SAGCN).
In this paper, we explore the fusion of the LLMs with conventional recommendation techniques to enhance both the accuracy and interoperability of recommendations. As shown in Fig. 2, our approach begins with the sentiment analysis of user-item reviews, utilizing LLMs with the chain-based prompting strategy to extract semantic aspect-aware interactions from each review. These extracted semantic aspect-aware reviews shed light on the fine-grained interactions underlying each user-item relationship. On the basis of these insights, we introduce a semantic aspect-based graph convolution network to enhance the user and item representations by leveraging these semantic aspect-aware interactions. To this end, our approach can properly extract and leverage the semantic aspect-aware interactions among each user-item pair for better user and item representation learning.
Next, we detail our method for semantic aspect-aware review extraction and semantic aspect-based graph convolution network.
3.3. Chain-based Prompting Strategy
In this section, we conduct sentiment analysis on user reviews using LLM to discern user-item interactions on various semantic aspects. Despite the powerful capability of LLMs in sentiment analysis, it is still a challenge to accurately identify the semantic aspect-aware interactions in each review without any instruction. This is because user reviews often contain complex, nuanced language that varies widely in terms of vocabulary, syntax, and semantics. This complexity makes it difficult to consistently identify specific aspects and associated sentiments. On the other side, while LLMs are adept at grasping context, the depth and breadth of their understanding can be limited by the training data. They might not always capture the full context or the subtle cues that indicate sentiment towards specific aspects, especially when these aspects are not explicitly mentioned. To alleviate these problems, as depicted in Fig. 3, we introduce a chain-based prompting strategy that begins by collecting the semantic aspects from all user reviews. Subsequently, we refine and select high-quality semantic aspects from all reviews and use these as the instruction information to generate the semantic aspect-aware reviews for each review. Finally, we identify aspect-aware interactions based on the generated reviews.
3.3.1. Semantic Aspects Extraction
Figure 3. Semantic Aspects and Semantic Aspect-aware Reviews Extraction.
The semantic aspects vary depending on the recommendation scenario. For instance, when buying dresses, style is crucial, but for office products, ease of use is the main concern. In this work, we first utilize the understanding capabilities of LLMs to extract unseen semantic aspects from the dataset. As illustrated in Fig. 3(a), the semantic aspects extraction prompt is designed as follows:
[…]
3.3.2. Semantic Aspect-Aware Review Extraction
Empirically, LLMs tend to generate more accurate results when provided with detailed prompts. To accurately identify the semantic aspects within each review, we further employ LLM to generate high-quality semantic aspect-aware reviews based on the obtained semantic aspect set $\mathcal{A}$. As shown in Fig. 1, an aspect-aware review provides a summary that identifies the presence of specific aspects within a review. Since we can focus on specific aspects, high-quality semantic aspect-aware reviews can be obtained. For instance, on the Baby Dataset, we get a semantic aspect set containing Quality, Comfort, Durability, Design, Functionality, Ease of use, Price, and Size. As illustrated in Fig. 3(b), we adapt the previous prompt to specific semantic aspects:
  • Prompt 2: A person bought a product and commented that ¡review¿. Tell me from which perspectives the customer gave this review, e.g., {$a^{1},a^{2},\cdots,a^{N}$}. Answer point by point.
The LLM identifies the semantic aspect-aware review for the aspects we input. Then, we identify the semantic aspect-aware interactions between the user and the item based on the generated semantic-aware reviews. For example, if the LLM generates the output: “The customer mentioned that the bins were not durable and that the fabric disintegrated when burned with a candle.”, it indicates an interaction on Functionality. On the other hand, if the LLM generates the output: “The customer did not mention anything about the ease of use of the bins.”, which means that there is no interaction in this aspect. Note that the time costs for the Chain-based Prompting strategy vary based on the length of input and output tokens. Based on our empirical evaluations, the process takes approximately 4 to 10 seconds per review.
3.4. Semantic Aspect-Based Graph Convolution Network
Figure 4. Overview of our SAGCN model.
Based on the extracted semantic-aware user-item interactions for semantic aspects, we seek to model users’ preferences on various aspects with GCN-based techniques. The structure of our SAGCN is illustrated in Fig. 4, and it is composed of three key components: 1) embedding initialization, which initialized node (users and items) embeddings for the subsequent learning in each aspect-specific interaction graph; 2) embedding propagation, in which the GCN technique (He et al., 2020) is adopted to learn the user and item embeddings by performing graph convolution on each semantic aspect-based graph; 3) embedding generation, which aggregates the embeddings learned from each aspect-specific graph and combines the embeddings of different aspects.
3.4.1. Semantic Aspect-based Graph Construction
@@ 3.4.2. Embedding Initialization
In the proposed SAGCN, the vectors of users and items are composed of embeddings from A aspects. Following existing GCN-based recommendation models (Wang et al., 2019b; He et al., 2020; Liu et al., 2021), we initialize the embedding vectors for each aspect. As an example, the ID embeddings of user $u\in\mathcal{U}$ for different aspects are initialized as:
@@
where $\bm{e_{u}^{(0)}}\in\mathds{R}^{d\times A}$ is ID embedding capturing intrinsic characteristics of user $u$; $\bm{e_{u}^{(a,0)}}\in\mathds{R}^{{d}}$ is $u$’s representation in aspect $a$. Analogously, $\bm{e_{i}}^{(0)}=\{\bm{e_{i}^{(1,0)},e_{i}^{(2,0)},\cdots,e_{i}^{(A,0)}}\}$ is defined as the ID embeddings of item $i\in\mathcal{I}$. In addition, it is worth noting that our model randomly and separately initializes the embeddings of the user IDs and item IDs for various aspects.
@@ 3.4.3. Embedding Propagation
To exploit the high-order connectives on each aspect to learn user and item representations, we propagate embeddings across various semantic aspect-based graphs. Take the semantic aspect-specific graph $\mathcal{G}_{a}$ as an example, after propagating embeddings through $k$ graph convolution layers, we define $\bm{e_{u}^{(a,k)}}$ and $\bm{e_{i}^{(a,k)}}$ to denote the embedding of user $u$ and item $i$, which are propagated from $k-$th order neighboring nodes, where $k\geq 0$, the embedding propagation is formulated as:
@@ 3.4.5. Model Training
In this work, we aim to recommend a list of items to the target users by computing their preference score. For the optimization, we utilize the pairwise learning method, which is similar to the approach employed in previous works (He et al., 2020; Liu et al., 2021). To train our SAGCN model, we construct a triplet of $\{u,i^{+},i^{-}\}$, comprising a positive item (an observed interaction between $u$ and $i^{+}$) and a negative item (an unobserved interaction between $u$ and $i^{-}$). The objective function is formulated as follows:
@@
where $\mathcal{O}$ represent the training set, defined as $\mathcal{O}=\{(u,i^{+},i^{-})|(u,i^{+})\in\mathcal{R^{+}},(u,i^{-})\in\mathcal{R^{-}}\}$. Here, $\mathcal{R^{+}}$ indicates the observed interactions in the training dataset, while $\mathcal{R^{-}}$ denotes a set of sampled unobserved interactions. Additionally, $\lambda$ represents the regularization weight, and $\Theta$ stands for the model parameters. We employ $L_{2}$ regularization to mitigate overfitting. To optimize our model and update its parameters, we utilize the mini-batch Adam technique, as described in (Kingma and Ba, 2015).
@@ 4.1. Experimental setup
To evaluate the effectiveness of our approach, we conducted experiments on Amazon Product Datasets and Goodreads Review Datasets.
@@
Amazon Product Datasets  [^2] http://jmcauley.ucsd.edu/data/amazon. These datasets stem from the extensive product review system of Amazon. It includes user-product interactions, reviews, and metadata. In our experiments, we focused solely on user-product interactions and user reviews. Three product categories from the 5-core version of these datasets are selected for our evaluation: Office Products (Office), Baby, and Clothing. This indicates that both users and items have a minimum of 5 interactions.
Goodreads Review Datasets  [^3] https://www.goodreads.com/. The datasets stem from the product review of GoodReads, which is the largest site for readers and book recommendations. It collected user-book interactions, reviews, and book metadata with anonymized user IDs and book IDs. In our experiments, we only use user-book interactions and review text.
In line with previous studies on conventional recommendation methods (He et al., 2020; Liu et al., 2021), we randomly split each user’s interaction into training and testing sets with a ratio of 80:20 for each dataset. Within the training set, we randomly select 10% of interactions to establish the validation set for hyper-parameter optimization. The pairwise learning strategy is adopted for parameter optimization. Specifically, the observed user-item interactions in the training set were treated as positive instances. For each positive instance, we paired it with a random sampling of negative instances that the user had not interacted with yet. For a fair comparison, we consistently adopt the negative sampling strategy across our methods and competitors. To further validate the effectiveness of our approach, we compare our proposed SAGCN with a pure LLM-based method, and the details of test dataset construction for the LLM-based method can be seen in Section 4.3.
4.1.2. Compared Baselines.
To demonstrate the effectiveness of our approach, we compared our SAGCN with several state-of-the-art models, which include CF-based recommendation methods without side information ($i.e.$, NeuMF, GCMC, LightGCN, DGCF, NCL) and multimodal recommendation methods ($i.e.$, MMGCN, GRCN, LATTICE, BM3). For fair comparisons, we only adopt reviews as the side information for MMGCN, GRCN, LATTICE, and BM3. In addition, we devised a LLM-based recommendation method for comparison, the details are provided in Section 4.3.
  • NeuMF (He et al., 2017): This is a state-of-the-art neural collaborative filtering method that present a general framework for collaborative filtering based on neural networks. It innovatively introduces multiple hidden layers to capture the non-linear interactions between the latent features of users and items.
  • GCMC (van den Berg et al., 2018): It is a graph-based auto-encoder framework for matrix completion. This method leverages the GCN technique (Kipf and Welling, 2017) on the user-item interaction graph and solely uses one convolutional layer to exploit direct connections between the nodes of users and items.
  • NGCF (Wang et al., 2019b): This is the very first GCN-based recommendation model that incorporates high-order connectivity of user-item interactions. It encodes the collaborative signal from high-order neighbors by performing embedding propagation on the user-item bipartite graph.
  • LightGCN (He et al., 2020): It is a state-of-the-art graph collaborative filtering method. This model can be considered a simplified version of NGCF (Wang et al., 2019b), where certain components such as the feature transformation and non-linear activation module are removed.
  • NCL (Lin et al., 2022): This is a model-agnostic contrastive learning framework, which explicitly incorporates the potential neighbors into contrastive pairs. It aims to capture the correlation between a node and its prototype, thereby refining the neural graph collaborative filtering techniques.
  • LATTICE (Zhang et al., 2021): This method uses side information to construct an item-item graph via a modality-aware structure learning layer and perform graph convolution operations to aggregate information from high-order neighbors to enhance the item representations.
  • MMGCN (Wei et al., 2019): The model learns user preferences and item characteristics by propagating modality embeddings on the user-item interaction graph.
  • GRCN (Wei et al., 2020): This GCN-based multimedia recommendation model designs a graph refining layer to adjust the structure of the interaction graph adaptively based on the status of model training for discovering and pruning potential false-positive edges.
  • BM3 (Zhou et al., 2023): This self-supervised framework uses a latent representation dropout mechanism to generate the target view of a user or an item. Moreover, it is trained without negative samples by jointly optimizing three multi-modal objectives.
  • DeepCoNN  (Zheng et al., 2017): This model jointly learns item properties and user behavior from review text by using Convolutional Neural Networks (CNNs) to encode the textual data.
  • NARRE (Chen et al., 2018): The model uses two parallel CNNs to learn the latent representation of reviews. Then, an attention mechanism is designed to estimate the contributions of different reviews for the item and user.
  • RGCL (Shuai et al., 2022): This model constructs a review-aware user-item graph with feature-enhanced edges derived from reviews. Each edge feature incorporates both the user-item rating and the semantic information extracted from the corresponding review.
  • LightGCNLDA (Blei et al., 2003): To validate the ability of LLMs in understanding reviews, we designed a baseline that combines LDA with LightGCN. LDA extracts topics from the dataset, and we calculate the similarity between each review and these topics. Using these similarities, we construct topic-specific graphs, where LightGCN learns user and item representations through graph convolutions.
  • LlamaRec (Yue et al., 2023): The model is a two-stage ranking-based recommendation. This method first retrieves a series of candidate items by small-scale recommenders learning based on the user interaction history. Then, the model adopts an LLM-based ranking model to reranking the order of candidate items.
  • RLMRec (Ren et al., 2024): The model designs an LLM-based generation framework that generates user and item profiles based on interaction data and textual information, such as product descriptions, user reviews, and product titles. Then, these generated profiles are encoded as semantic information of LLMs and incorporated into the CF-based recommendation system.
4.1.3. Evaluation Metrics
Two widely-used evaluation metrics are employed for evaluating top-$K$ recommendations: Recall and Normalized Discounted Cumulative Gain (NDCG) (He et al., 2015). For conventional recommendation methods, the recommendation accuracy is calculated for each metric based on the top 10 and 20 results. The reported results are the average values across all testing users. In contrast, for the LLM-based method we have devised for comparison, we calculate the recommendation accuracy based on the top 1, 3, 5, and 7 results. In addition, the reported results are the average values across the sampled users. Note that we sample several users for evaluation due to the high resource and time consumption of LLMs.
4.1.4. Parameter Settings
The PyTorch framework  [^4] https://pytorch.org. is adopted to implement the proposed model. We employed Vicuna-13B   [^5] https://huggingface.co/lmsys/vicuna-13b-v1.5 to analyze the user reviews. It has been fine-tuned from LLaMA using supervised instruction-based fine-tuning and exhibits strong performance in sentiment analysis. In our experiments, all hyperparameters are carefully tuned. The user and item embedding size for all competitor methods is set to 64 across all datasets. The number of semantic aspects $N$ is defined as 8 across all four datasets. The embedding size associated with each aspect is also set to 64 within our proposed SAGCN. The mini-batch size is fixed to 1024. The learning rate for the optimizer is searched from $\{1e^{-5},1e^{-4},\cdots,1e^{1}\}$, and the model weight decay is searched in the range $\{1e^{-5},1e^{-4},\cdots,1e^{-1}\}$. Besides, the early stopping strategy is implemented to enhance training efficiency. Specifically, the training process will stop if recall@10 does not increase for 20 successive epochs.
4.2. Performance Comparison With Baseline Methods
Table 2. Performance of our method and the competitors over four datasets.
Datasets Office Baby Clothing Goodreads
Metrics R@10 N@10 R@20 N@20 R@10 N@10 R@20 N@20 R@10 N@10 R@20 N@20 R@10 N@10 R@20 N@20
NeuMF 5.14 3.89 8.12 5.21 3.11 2.11 4.85 2.69 0.94 0.54 1.50 0.71 9.24 7.45 14.63 9.02
GCMC 6.72 5.27 10.27 6.79 4.55 2.99 7.24 3.89 3.17 1.86 4.86 2.35 16.32 10.58 22.32 12.32
LightGCN 9.87 6.04 14.47 7.43 5.94 3.30 9.25 4.20 4.45 2.43 6.44 2.95 16.99 11.02 23.35 13.01
DGCF 9.95 6.25 14.37 7.52 5.90 3.27 9.20 4.14 4.67 2.68 6.91 3.32 16.89 11.22 21.76 13.12
IMP-GCN 10.11 6.36 14.47 7.71 6.24 3.49 9.56 4.38 4.80 2.76 7.11 3.40 18.05 12.07 23.79 14.12
NCL 10.07 6.30 14.40 7.65 6.15 3.42 9.45 4.30 4.76 2.74 7.10 3.37 17.69 11.64 23.55 13.68
MMGCN 5.74 3.42 9.39 4.54 3.95 2.17 6.46 2.85 2.42 1.29 3.76 1.64 11.24 7.85 15.33 9.48
GRCN 10.38 6.34 15.33 7.81 5.57 3.03 8.49 3.83 4.47 2.35 6.70 2.94 16.62 10.79 22.64 12.67
LATTICE 10.00 6.09 14.99 7.57 6.06 3.40 9.29 4.27 5.03 2.79 7.28 3.37 17.92 11.35 23.60 13.27
BM3 9.80 6.09 14.02 7.36 6.44 3.65 9.52 4.48 5.28 2.94 7.75 3.58 18.28 12.12 24.68 14.15
DeepCoNN 5.32 4.01 8.35 5.33 3.20 2.02 5.05 2.71 1.89 1.01 2.98 1.35 11.89 7.22 5.66 8.44
NARRE 6.12 4.78 9.41 6.15 4.02 2.32 6.14 2.98 2.37 1.32 3.62 1.69 12.79 9.32 16.84 10.89
RGCL 7.89 5.69 12.40 7.02 5.22 2.54 8.20 3.19 3.32 1.93 5.22 2.41 16.55 10.43 21.79 12.19
LightGCNLDA 10.12 6.24 15.12 7.67 5.98 3.44 9.41 4.29 5.17 2.88 7.50 3.44 16.73 10.87 22.07 12.70
LlamaRec 9.89 5.83 14.44 7.20 5.62 3.11 8.83 3.85 3.72 2.17 5.45 2.68 16.67 10.82 21.80 12.66
RLMRec 10.1 6.35 14.87 7.78 5.84 3.37 8.91 4.30 4.48 2.60 6.65 3.08 16.72 11.02 22.02 12.71
SAGCN 11.71* 7.34* 16.71* 8.84* 7.35* 4.23* 10.56* 5.09* 6.07* 3.58* 8.44* 4.20* 19.40* 13.15* 26.17* 15.14*
Imporv. 12.81% 15.59% 9.00% 13.19% 14.13% 15.89% 10.46% 13.61% 14.96% 21.77% 8.90% 17.32% 6.15% 8.50% 6.03% 7.03%